OpenStudy (anonymous):

A baseball player hits a ball toward the outfield. The height h of the ball in feet is modeled by h(t) = -16t2 + 22t + 3, where t is the time in seconds. In addition, the function d(t) = 85t models the horizontal distance d traveled by the ball. If no one catches the ball, how long will it stay in the air?

OpenStudy (anonymous):

put 0 in for the height to get how long its in the air -16t^2+22t+3=0 now solve for t

OpenStudy (anonymous):

I'm getting a really weird fraction. somehow before, i got 3 as my answer. am i even close?

OpenStudy (anonymous):

so i got 0.09, and if i have to round, would it be 0.1?

OpenStudy (anonymous):

factors like (2t-3)(8t+1) 8t+1=0 8t=-1 t=-1/8 throw that one away because its negative 2t-3=0 2t=3 t=3/2 or 1.5 seconds

OpenStudy (anonymous):

ohhh okay! i get it now. thank you so much!